# On the Complexity of Minimizing Energy Consumption of Partitioning DAG Tasks

### v1

@@ Abstract
We study a graph partition problem where we are given a directed acyclic graph (DAG) whose vertices and arcs can be respectively regarded as tasks and dependencies among tasks. The objective of the problem is to minimize the total energy consumed for completing these tasks by assigning the tasks to $k$ heterogeneous machines. We first show that the problem is NP-hard. Then, we present polynomial-time algorithms for two special cases where there are only two machines and where the input DAG is a directed path. Finally, we study a natural variant where there are only two machines with one of them being capable of executing a limited number of tasks. We show that this special case remains computationally hard.
keywords:
graph partition problems , NP-hard , DAG , task allocation problems
journal: Theoretical Computer Science
[…]
@@ 1 Introduction
Tasks that are represented by directed acyclic graphs (DAG tasks) are ubiquitous in many applications, including for instance cloud computing, deep neural network, workflow scheduling, etc. [1, 2, 3, 4, 5, 6]. In this paper, we investigate the complexity of a new graph partition problem which models the scenario where DAG tasks are deemed to be assigned to $k$ heterogeneous machines (e.g., execution units in distributed systems, clusters of cores in heterogeneous multicore systems, etc.), with the objective to minimize the energy consumption for the computation of these tasks under natural restrictions. More precisely, in this problem, we are given a DAG whose vertices represent tasks, in which the energy consumption of a task depends on which machine is allocated for its execution. An arc from a task $a$ to a task $b$ means that the computation of $b$ requires the output of task $a$. So, if $a$ and $b$ are assigned to different machines, the output of $a$ needs to be transferred to the machine executing the task $b$, which also incurs energy consumption. We note that when there are multiple outneighbors of a task $a$ assigned to a machine $i$ different from that of $a$, we need only to transfer once the output of $a$ to the machine $i$.
We assume that the energy consumption associated with the transfer primarily depends on the volume of transferred data and the impact of other factors is neglectable. Under this assumption, we may simply use a univariate function mapping tasks to numbers to define energy consumption for data transformation. A more general case is when, in addition to the volume of the data, energy consumption also takes into account the types of data, the identities of machines between which the data is transferred, etc. We prove that even in our special case, the problem is already NP-hard.
@@
A directed graph (digraph) is a tuple $G=(V,A)$ where $V$ is a set of vertices and $A$ is a set of arcs. Each arc is defined as an ordered pair of vertices. An arc from a vertex $v$ to a vertex $u$ is denoted by $(v,u)$. We say that the arc $(v,u)$ leaves $v$ and enters $u$. We also use $A(G)$ to denote the set of arcs of $G$, and use $V(G)$ to denote the set of vertices of $G$. The set of outneighbors (respectively, inneighbors) of a vertex $v\in V(G)$ is defined as $\Gamma^{+}_{G}{(v)}=\{u\in V(G):(v,u)\in A(G)\}$ (respectively, $\Gamma^{-}_{G}{(v)}=\{u\in V(G):(u,v)\in A(G)\}$). The outdegree (respectively, indegree) of $v$ is the number of outneighbors (respectively, inneighbors) of $v$ in $G$. The set of closed outneighbors (respectively, closed inneighbors) of $v\in V(G)$ is defined as $\Gamma^{+}_{G}{[v]}=\Gamma^{+}_{G}{(v)}\cup\{v\}$ (respectively, $\Gamma^{-}_{G}{[v]}=\Gamma^{-}_{G}{(v)}\cup\{v\}$). Vertices of outdegree zero are called sinks, and those of indegree zero are called sources of $G$. A DAG is a digraph without directed cycles.
For $v\in V(G)$ and $S\subseteq V(G)$, let $\Gamma^{+}_{G}{(v,S)}=\Gamma^{+}_{G}{(v)}\cap S$. In addition, we define
${\bf{1}}_{G}(v,S)=\begin{cases}1,&\text{if }\Gamma^{+}_{G}{(v,S)}\neq\emptyset,\\ 0,&\text{otherwise}.\\ \end{cases}$
For an integer $i$, we use $[i]$ to denote the set of all positive integers no greater than $i$. For a function $f:X\rightarrow Y$, we use $f^{-1}(y)$ to denote the set consisting of all $x\in X$ such that $f(x)=y$. For two functions $f:V(G)\rightarrow[k]$ and $q:V(G)\rightarrow\mathbb{R}_{\geq 0}$, we define
@@
In Equality (1), $f$ is a function assigning vertices of $G$ to $k$ machines, $q$ defines the energy consumption for transferring the outputs of vertices in $G$, and ${\textsf{Tran}_{G}(f,q)}$ is exactly the total amount of energy consumption for transferring all necessary data under the assignment $f$.
Energy-Saving Partition of DAG (ESP-DAG)
@@
Question: Is there an assignment function $f:V(G)\rightarrow[k]$ so that $\left(\sum_{v\in V(G)}p(v,f(v))\right)+{\textsf{Tran}_{G}(f,q)}\leq r?$
In the formulation, $p(v,i)$ for each $v\in V(G)$ and each $i\in[k]$ is the amount of energy consumed for the computation of task $v$ in the machine $i$.
We also study a natural variant of the special case where $k=2$ under the restriction that one of the two machines is capable of executing a limited number of tasks. For simplicity of presentation, we reformulate the special case as follows. For a graph $G$, two disjoint subsets $V_{1},V_{2}\subseteq V(G)$, and a function $q:V(G)\rightarrow\mathbb{R}$, let
@@
We first establish a complexity dichotomy for ESP-DAG with respect to the number of machines: the problem is NP-hard if there are at least three machines (Theorem 1), but becomes polynomial-time solvable when there are two machines (Theorem 2). Afterwards, we show that when the input DAG degenerates to a directed path [^1] Such a DAG is also called a chain in the literature. , the problem admits a pseudo polynomial-time algorithm, regardless of the number of machines (Theorem 3). In contrast to ESP-DAG, we show that the size-bounded variant SB-ESBP-DAG, where there are only two machines, is computationally hard: it is W[1]-hard with respect to the parameter $\ell$ (Theorem 6). As a byproduct of this result, we show that a variant of the minimum cut problem is W[1]-hard with respect to a natural parameter (Theorem 4), strengthening its NP-hardness studied in [11].
@@ 2 Related Works
To the best of our knowledge, ESP-DAG has not been studied from a theoretical point of view in the literature. However, the problem and its variant defined above belong to the gigantic family of graph partition problems which have been extensively and intensively studied in the literature. These problems aim at dividing either the vertices or the arcs of a given graph (undirected or directed) into several sets so that certain structural properties are met or certain optimization objections are achieved (see, e.g., [12, 13, 14, 15, 16, 17, 18, 19]).
More specifically, our problem falls into the category of task allocation problems [20, 21, 22, 23]. In the realm of task allocation problems, partitioning a DAG can achieve various objectives. These objectives include minimizing makespan, reducing energy consumption, decreasing communication costs, achieving load balancing, ensuring fault tolerance, etc. Our model aims to minimize energy consumption. In most of the previous works tackling DAG tasks, the cost of data transformation are arc-wisely defined: if a task $v$ assigned to a machine $i$ has multiple outneighbors assigned to a different machine $j$, the output of $v$ needs to be transferred multiple times from machine $i$ to machine $j$, one for each of $v$’s outneighbors assigned to the machine $j$. We build upon previous approaches by considering a more comprehensive energy consumption model for data transformation. We simplify multiple transfers of output data from one machine to another when a task has multiple outneighbors assigned to different machines. This refined modeling approach enables us to evaluate and minimize energy consumption more accurately in DAG-based task allocation.
One noticeable related problem where a similar energy model as ours is adopted is a one studied by Hu et al. [24] in 2019. Particularly, this problem takes as the same input as SB-ESBP-DAG with $\ell$ being dropped, and the problem consists in dividing $V(G)$ into two sets $V_{1}$ and $V_{2}$ to minimize $\sum_{v\in V_{i},i\in[2]}p(v,i)+\sum_{v\in V_{1}}q(v)\cdot{\bf{1}}_{G}(v,V_{2})$, under the restriction that there are no arcs from $V_{2}$ to $V_{1}$. For the problem, Hu et al. [24] proposed a polynomial-time algorithm which is, however, pointed out to be flawed by Li et al. [25].
@@ 3 Problems to Establish Our Results
Our results are obtained based on the the following problems.
An undirected graph $G=(V,E)$ is a tuple where $V$ is a set of vertices and $E$ is a set of edges. The set of vertices and the set of edges of $G$ are also denoted by $V(G)$ and $E(G)$, respectively. An edge between two vertices $v$ and $u$ is denoted by $\{v,u\}$.
For a function $w:S\rightarrow\mathbb{R}$ and a subset $S^{\prime}\subseteq S$, we define $w(S^{\prime})=\sum_{s\in S^{\prime}}w(s)$. For a graph $G$ and a subset $E^{\prime}\subseteq E(G)$, we use $G-E^{\prime}$ to denote the graph obtained from $G$ by removing all edges in $E^{\prime}$.
@@
Input: An undirected graph $G$, a weight function $w:E(G)\rightarrow\mathbb{R}_{\geq 0}$, a set $T$ of $k$ distinct vertices from $V(G)$ called terminals, and a number $r$.
Question: Is there a subset $E^{\prime}\subseteq E(G)$ such that $w(E^{\prime})\leq r$ and vertices from $T$ are pairwise disconnected in $G-E^{\prime}$?
Equivalently, Multiway Cut determines if we can partition $V(G)$ into $k$ sets, each containing exactly one terminal, so that the total weight of edges crossing the partition does not exceed $r$. It is known that Multiway Cut is NP-hard for every $k\geq 3$, but becomes polynomial-time solvable if $k=2$ [29]. (Multiway Cut with two terminals is exactly the decision version of the classic problem Minimum $s$-$t$-Cut.)
The problem Size Bounded Minimum $s$-$t$-Cut (SBM-$s$-$t$-Cut) takes as the same input as Multiway Cut where $k=2$ together with an integer $\ell$, and determines if there is a bipartition $(V_{s},V_{t})$ of $V(G)$ so that $|V_{s}|\leq\ell$, $s\in V_{s}$, $t\in V_{t}$, and the total weight of edges between $V_{s}$ and $V_{t}$ in $G$ is at most $r$. SBM-$s$-$t$-Cut is NP-hard [11].
A clique in an undirected graph is a subset of pairwise adjacent vertices.
@@
Input: An undirected graph $G$ and an integer $\ell$.
@@
Clique is a well-known NP-hard problem [30]. Moreover, it is W[1]-hard with respect to $\kappa$ even when restricted to regular graphs [31, 32, 33].
For a digraph $G$ and a subset $A^{\prime}$ of arcs, $G-A^{\prime}$ denotes the digraph obtained from $G$ by removing all arcs in $A^{\prime}$. For two disjoint subsets $X,Y\subseteq V(G)$, let ${{\textsf{cut}}}_{G}(X,Y)$ be the set of all arcs from $X$ to $Y$, i.e., ${{\textsf{cut}}}_{G}(X,Y)=\{(v,u)\in A(G):v\in X,u\in Y\}$.
Directed Minimum $s$-$t$ Cut (DM-$s$-$t$-Cut)
Input: A digraph $G$, a weight function $w:A(G)\rightarrow\mathbb{R}_{\geq 0}$, two distinct vertices $s$ and $t$ in $G$, an integer $r$.
Question: Is there a subset $A^{\prime}\subseteq A(G)$ so that $w(A^{\prime})\leq r$ and there is no directed path from $s$ to $t$ in $G-A^{\prime}$?
Equivalently, DM-$s$-$t$-Cut determines if we can partition $V(G)$ into two sets $V_{s}$ and $V_{t}$ so that $s\in V_{s}$, $t\in V_{t}$, and $w({{\textsf{cut}}}_{G}(V_{s},V_{t}))\leq r$. It is known that DM-$s$-$t$-Cut is polynomial-time solvable [16]).
@@ 4 Energy-Saving Partition of DAG
We prove Theorem 1 by a reduction from Multiway Cut to ESP-DAG. Let $(G,w,T,r)$ be an instance of Multiway Cut, where $T=\{t_{1},t_{2},\dots,t_{k}\}$. Without loss of generality, assume that $k\geq 2$. To construct an instance of ESP-DAG, we first arbitrarily orient $G$ into a DAG (e.g., we arbitrarily fix a linear order on $V(G)$ and orient edges forwardly). Let $\overrightarrow{G}$ denote this DAG. It holds that $V(G)=V(\overrightarrow{G})$. Then, we construct a graph $G^{\prime}$ obtained from $\overrightarrow{G}$ by subdividing all arcs: for each arc $(v,u)\in A(\overrightarrow{G})$, we introduce one new vertex $a(v,u)$, add the arcs $(v,a(v,u))$ and $(a(v,u),u)$, and remove the arc $(v,u)$. Clearly, each newly introduced vertex $a(v,u)$ has exactly one outneighbor $u$ and has exactly one inneighbor $v$. Moreover, $G^{\prime}$ remains as a DAG.
Now we construct two functions $p:V(G^{\prime})\times[k]\rightarrow\mathbb{R}_{\geq 0}$ and $q:V(G^{\prime})\rightarrow\mathbb{R}_{\geq 0}$ as follows. For each $x\in V(G^{\prime})$, we define
$p(x,i)=\begin{cases}+\infty,&\text{if }x=t_{j},j\in[k]\setminus\{i\},\\ 0,&\text{otherwise}.\\ \end{cases}$ (3)
For each $v\in V(\overrightarrow{G})$, we define $q(v)=+\infty$, and for each newly introduced vertex $a(v,u)\in V(G^{\prime})\setminus V(\overrightarrow{G})$ where $(v,u)\in A(\overrightarrow{G})$, we define $q(a(v,u))=w(\{v,u\})$.
The instance of ESP-DAG is $(G^{\prime},p,q,r)$. The reduction clearly runs in polynomial time. In the following, we show its correctness. For each $v\in V(G)$, let

### v3

@@ Abstract
We study a graph partition problem where the input is a directed acyclic graph (DAG) representing tasks as vertices and dependencies between tasks as arcs. The goal is to assign the tasks to $k$ heterogeneous machines in a way that minimizes the total energy consumed for completing the tasks. We first show that the problem is NP-hard. Then, we present polynomial-time algorithms for two special cases: one where there are only two machines, and another where the input DAG is a directed path. Finally, we examine a variant where there are only two machines, with one capable of executing a limited number of tasks, and demonstrate that this special case remains computationally hard.
@@ 1 Introduction
Tasks that are represented by directed acyclic graphs (DAG tasks) are ubiquitous in many applications, including for instance cloud computing, deep neural network, workflow scheduling, etc. [16, 23, 24, 26, 33, 35]. In this paper, we investigate the complexity of a new graph partition problem which models the scenario where DAG tasks are deemed to be assigned to $k$ heterogeneous machines (e.g., execution units in distributed systems, clusters of cores in heterogeneous multicore systems, etc.), with the objective to minimize the energy consumption for the computation of these tasks under natural restrictions. More precisely, in this problem, we are given a DAG whose vertices represent tasks, in which the energy consumption of a task depends on which machine is allocated for its execution. An arc from a task $a$ to a task $b$ means that the computation of $b$ requires the output of task $a$. So, if $a$ and $b$ are assigned to different machines, the output of $a$ needs to be transferred to the machine executing the task $b$, which also incurs energy consumption. We note that when there are multiple outneighbors of a task $a$ assigned to a machine $i$ different from that of $a$, we need to transfer the output of $a$ to the machine $i$ only once.
We assume that the energy consumption associated with data transfer depends solely on the volume of transferred data, and that the impact of other factors is negligible. Under this assumption, energy consumption for data transfer can be modeled as a univariate function that maps tasks to numerical values. In a more general scenario, energy consumption could also consider additional factors such as the type of data being transferred, the identities of the machines involved, and other relevant parameters. Nevertheless, we prove that even in our simplified case, the problem remains NP-hard.
@@
A directed graph (digraph) is a tuple $G=(V,A)$ where $V$ is a set of vertices and $A$ is a set of arcs. Each arc is defined as an ordered pair of vertices. An arc from a vertex $v$ to a vertex $u$ is denoted by $(v,u)$. We say that the arc $(v,u)$ leaves $v$ and enters $u$. We also use $A(G)$ to denote the set of arcs of $G$, and use $V(G)$ to denote the set of vertices of $G$. The set of outneighbors (respectively, inneighbors) of a vertex $v\in V(G)$ is defined as $N^{+}_{G}{(v)}=\{u\in V(G):(v,u)\in A(G)\}$ (respectively, $N^{-}_{G}{(v)}=\{u\in V(G):(u,v)\in A(G)\}$). The set of closed outneighbors of $v\in V(G)$ is defined as $N^{+}_{G}{[v]}=N^{+}_{G}{(v)}\cup\{v\}$. Vertices without any outneighbors are called sinks, and those without inneighbors are called sources of $G$. A DAG is a digraph without directed cycles. For a vertex $v\in V(G)$ and a subset $S\subseteq V(G)$, we define
${\bf{1}}_{G}(v,S)=\begin{cases}1,&\text{if }N^{+}_{G}{(v)}\cap S\neq\emptyset,\\ 0,&\text{otherwise},\\ \end{cases}$
which indicates whether $v$ has at least one outneighbor from $S$ in $G$.
For a function $f:X\rightarrow Y$ and an element $y\in Y$, we use $f^{-1}(y)$ to denote the set consisting of all $x\in X$ such that $f(x)=y$.
For an integer $k$, let $[k]$ denote the set of all positive integers less than or equal to $k$. We also use $[k]$ to represent a set of $k$ machines. Let $f:V(G)\rightarrow[k]$ be a function that assigns the vertices of $G$ to the $k$ machines. Additionally, let $p:V(G)\times[k]\rightarrow\mathbb{R}_{\geq 0}$ be a function where $p(v,i)$ defines the energy consumed for computing task $v$ on machine $i$, for each $v\in V(G)$ and $i\in[k]$. Furthermore, let $q:V(G)\rightarrow\mathbb{R}_{\geq 0}$ be a function that specifies the energy consumption for transferring the outputs of the vertices in $G$. The total energy consumption required to transfer all necessary data under the assignment $f$ is then defined as:
@@
The Energy-Saving Partition of DAG (ESP-DAG) problem is defined as follows.
ESP-DAG
@@
Question: Is there an assignment $f:V(G)\rightarrow[k]$ so that $\left(\sum_{v\in V(G)}p(v,f(v))\right)+{\textsf{Tran}_{G}(f,q)}\leq r?$
See Figure 1 for an illustration of the problem.
Figure 1: An illustration of the ESP-DAG problem. (a) A DAG $G$. The number next to each task (vertex) represents the energy consumption required to transfer the task’s output. (b) Energy consumption for executing the tasks on three machines: M1, M2, and M3. (c) and (d) are two different assignments of the tasks to the machines. The corresponding energy consumption are respectively $(1+8+2+2+4)+(2+3)=22$ (c) and $(1+1+2+2+4)+(2+3)=15$ (d).
We also investigate a natural variant of the special case where $k=2$, with the restriction that one of the two machines can execute only a limited number of tasks. For clarity, we reformulate this variant as follows. For a digraph $G$, two disjoint subsets $V_{1},V_{2}\subseteq V(G)$, and a function $q:V(G)\rightarrow\mathbb{R}_{\geq 0}$, let
@@
We first establish a complexity dichotomy for ESP-DAG with respect to the number of machines: the problem is NP-hard if there are at least three machines (Theorem 1), but becomes polynomial-time solvable when there are two machines (Theorem 2). Afterwards, we show that when the input DAG degenerates to a directed path [^1] Such a DAG is also called a chain in the literature., the problem becomes polynomial-time solvable, regardless of the number of machines (Theorem 3). Additionally, we show that the size-bounded variant SB-ESBP-DAG, where there are only two machines, is computationally hard: it is W[1]-hard with respect to the parameter $\ell$ (Theorem 5). As a byproduct of this result, we show that a variant of a minimum cut problem is W[1]-hard with respect to a natural parameter (Theorem 4), strengthening its NP-hardness studied in [9].
@@ 2 Related Works
To the best of our knowledge, the ESP-DAG problem has not been studied in the literature. However, the problem and its variant defined in the previous section belong to the gigantic family of graph partition problems, which have been extensively and intensively studied in the literature. These problems aim at dividing either the vertices or the arcs/edges of a given digraph/graph into several sets so that certain structural properties are met or certain optimization objections are achieved (see, e.g., [2, 3, 4, 6, 7, 13, 31, 34]).
More specifically, the ESP-DAG problem falls into the category of task allocation problems [17, 22, 27, 32]. Partitioning a DAG in such problems can achieve various objectives, including minimizing makespan, reducing energy consumption, decreasing communication costs, achieving load balancing, and ensuring fault tolerance. Our model aims to minimize energy consumption. In most of the previous works tackling DAG tasks, the cost of data transformation are arc-wisely defined: if a task $v$ assigned to a machine $i$ has multiple outneighbors assigned to a different machine $j$, the output of $v$ needs to be transferred multiple times from machine $i$ to machine $j$, one for each of $v$’s outneighbors assigned to machine $j$. We build upon previous approaches by considering a more comprehensive energy consumption model for data transformation. We simplify multiple transfers of output data from one machine to another when a task has multiple outneighbors assigned to different machines. This refined modeling approach enables us to evaluate and minimize energy consumption more accurately in DAG-based task allocation.
One noticeable related problem where a similar energy model as ours is adopted is a one studied by Hu et al. [19] in 2019. Particularly, this problem takes as the same input as SB-ESBP-DAG with $\ell$ being dropped, and the problem consists in dividing $V(G)$ into two disjoint sets $V_{1}$ and $V_{2}$ to minimize $\sum_{v\in V_{i},i\in[2]}p(v,i)+\sum_{v\in V_{1}}q(v)\cdot{\bf{1}}_{G}(v,V_{2})$, under the restriction that there are no arcs from $V_{2}$ to $V_{1}$. For the problem, Hu et al. [19] proposed a polynomial-time algorithm which is, however, pointed out to be flawed by Li et al. [25].
@@ 3 Problems to Establish Our Results
Our results are obtained based on the following problems.
An undirected graph $G=(V,E)$ is a tuple where $V$ is a set of vertices and $E$ is a set of edges. For simplicity, throughout the remainder of this paper, we will refer to undirected graphs simply as graphs. The set of vertices and the set of edges of $G$ are also denoted by $V(G)$ and $E(G)$, respectively. An edge between two vertices $v$ and $u$ is denoted by $\{v,u\}$. Two vertices are connected in $G$ if there is path between them. For a subset $E^{\prime}\subseteq E(G)$, we use $G-E^{\prime}$ to denote the graph obtained from $G$ by removing all edges in $E^{\prime}$.
For a function $w:S\rightarrow\mathbb{R}$ and a subset $S^{\prime}\subseteq S$, let $w(S^{\prime})=\sum_{s\in S^{\prime}}w(s)$.
@@
Input: A graph $G$, a weight function $w:E(G)\rightarrow\mathbb{R}_{\geq 0}$, a set $T=\{t_{1},t_{2},\dots,t_{k}\}$ of $k$ distinct vertices in $G$ called terminals, and a number $r$.
Question: Is there a subset $E^{\prime}\subseteq E(G)$ such that $w(E^{\prime})\leq r$, and every two distinct $t_{i},t_{j}\in T$ are disconnected in $G-E^{\prime}$?
Equivalently, Multiway Cut determines if there is a partition $(V_{i})_{i\in[k]}$ of $V(G)$ such that $t_{i}\in V_{i}$ for all $i\in[k]$ and the total weight of edges crossing the partition does not exceed $r$. The Multiway Cut problem is NP-hard for every $k\geq 3$, but becomes polynomial-time solvable if $k=2$ [12]. Multiway Cut with two terminals is exactly the decision version of the classic problem Minimum $s$-$t$-Cut. The Size Bounded Minimum $s$-$t$-Cut problem (SBM-$s$-$t$-Cut) takes as input a graph $G$, a weight function $w:E(G)\rightarrow\mathbb{R}_{\geq 0}$, a pair $\{s,t\}\subseteq V(G)$ of vertices, and two numbers $r$ and $\ell$. The problem asks whether there exists a bipartition $(V_{s},V_{t})$ of $V(G)$ such that $|V_{s}|\leq\ell$, $s\in V_{s}$, $t\in V_{t}$, and the total weight of the edges between $V_{s}$ and $V_{t}$ in $G$ is at most $r$. It is known that the SBM-$s$-$t$-Cut problem is NP-hard [9].
A clique in a graph is a subset of pairwise adjacent vertices.
@@
Input: A graph $G$ and an integer $\ell$.
@@
Clique is a well-known NP-hard problem [21]. Moreover, it is W[1]-hard with respect to $\ell$ even when restricted to regular graphs [8, 28, 29].
For a digraph $G$ and a subset $A^{\prime}$ of arcs, $G-A^{\prime}$ denotes the digraph obtained from $G$ by removing all arcs in $A^{\prime}$. For two disjoint subsets $X,Y\subseteq V(G)$, let ${{\textsf{cut}}}_{G}(X,Y)$ be the set of all arcs from $X$ to $Y$, i.e., ${{\textsf{cut}}}_{G}(X,Y)=\{(v,u)\in A(G):v\in X,u\in Y\}$. The decision version of the Directed Minimum $s$-$t$ Cut problem (DM-$s$-$t$-Cut) is defined as follows.
DM-$s$-$t$-Cut
Input: A digraph $G$, a weight function $w:A(G)\rightarrow\mathbb{R}_{\geq 0}$, a pair $\{s,t\}\subseteq V(G)$ of two vertices, an integer $r$.
Question: Is there a bipartition $(V_{s},V_{t})$ of $V(G)$ such that $s\in V_{s}$, $t\in V_{t}$, and $w({{\textsf{cut}}}_{G}(V_{s},V_{t}))\leq r$?
Equivalently, DM-$s$-$t$-Cut determines if there is a subset of arcs $A^{\prime}\subseteq A(G)$ such that $w(A^{\prime})\leq r$ and there is no directed path from $s$ to $t$ in $G-A^{\prime}$. We call $w({{\textsf{cut}}}_{G}(V_{s},V_{t}))$ the size of ${{\textsf{cut}}}_{G}(V_{s},V_{t})$ with respect to $w$. The DM-$s$-$t$-Cut problem is polynomial-time solvable [13].
@@ 4 Energy-Saving Partition of DAG
We prove Theorem 1 by a reduction from the Multiway Cut problem to the ESP-DAG problem. Let $(G,w,T,r)$ be an instance of the Multiway Cut problem, where $T=\{t_{1},t_{2},\dots,t_{k}\}$. Without loss of generality, we assume that $k\geq 3$. To construct an instance of the ESP-DAG problem, we first arbitrarily fix a linear order on $V(G)$ and orient edges of $G$ forwardly. Let $\overrightarrow{G}$ denote the resulting graph, which is clearly a DAG. It holds that $V(G)=V(\overrightarrow{G})$. Then, we construct a digraph $G^{\prime}$ obtained from $\overrightarrow{G}$ by subdividing all arcs: for each arc $(v,u)\in A(\overrightarrow{G})$, we introduce one new vertex $a(v,u)$, add the arcs $(v,a(v,u))$ and $(a(v,u),u)$, and remove the arc $(v,u)$. Clearly, each newly introduced vertex $a(v,u)$ has exactly one outneighbor $u$ and exactly one inneighbor $v$. Moreover, $G^{\prime}$ remains as a DAG. See Figure 2 for an illustration.
Figure 2: An illustration of the reduction in the proof of Theorem 1. The digraph $\overrightarrow{G}$ is constructed based on the linear ordering $(t_{1},x,y,z,t_{2},t_{3})$. Dark vertices in $G^{\prime}$ are newly introduced vertices in the reduction.
Now we construct two functions $p:V(G^{\prime})\times[k]\rightarrow\mathbb{R}_{\geq 0}$ and $q:V(G^{\prime})\rightarrow\mathbb{R}_{\geq 0}$ as follows. For each $v\in V(G^{\prime})$, we define
$p(v,i)=\begin{cases}+\infty,&\text{if }v=t_{j},j\in[k]\setminus\{i\},\\ 0,&\text{otherwise}.\\ \end{cases}$ (3)
For each $v\in V(\overrightarrow{G})$, we define $q(v)=+\infty$, and for each newly introduced vertex $a(v,u)\in V(G^{\prime})\setminus V(\overrightarrow{G})$ where $(v,u)\in A(\overrightarrow{G})$, we define $q(a(v,u))=w(\{v,u\})$. An instance $(G^{\prime},p,q,r)$ of the ESP-DAG problem is constructed.
The above reduction clearly runs in polynomial time. In the following, we show its correctness. For each $v\in V(G)$, let